Plot synopsis of Jaws

A group of so-called government funded ‘experts’ whip up alarmist fears of a killer shark off the coast of Amity, a sea side town. Their goal is to destroy the local tourist industry, send Amity back to the dark ages and thus achieve their underlying socialist agenda of wealth redistribution. The heroes of this tale are the local major and business leaders who lead a successful audit of the alarmist claims and by doing so manage to delay action long enough that the beach remains open. In the end it turns out a shark has been eating people.

[From http://denialdepot.blogspot.com/2010/11/jaws-movie-review.html]